OpenStudy (anonymous):

The change in water vapor in a cloud is modeled by a polynomial function, C(x). Describe how to find the x-intercepts of C(x) and how to construct a rough graph of C(x) so that the meteorologist can predict when there will be no change in the water vapor. You may create a sample polynomial to be used in your explanations. @ganeshie8

OpenStudy (anonymous):

To find the x-intercept of C(x), we start by setting the equation ( C(x) ) to be equal to zero and then we solve for x. This can also be applied to any polynomial to find the x-intercept. This is what I have so far. I don't get the how to construct a rough graph part

ganeshie8 (ganeshie8):

Thats a very good start!

OpenStudy (anonymous):

Thank you!! ;v; I just don't know how to finish it off haha @ganeshie8

ganeshie8 (ganeshie8):

Next, cookup an example polynomial like : C(x) = x^3-x

OpenStudy (anonymous):

I can use that one right? @ganeshie8

ganeshie8 (ganeshie8):

set that equal to 0 and solve for x intercepts

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

x^3 - x = 0 can you solve x ?

OpenStudy (anonymous):

x= \[\sqrt[3]{x}\] ? @ganeshie8

ganeshie8 (ganeshie8):

not exactly x^3 - x = 0 factor x x(x^2-1) = 0

ganeshie8 (ganeshie8):

x = 0 or x^2 - 1 = 0 x^2 - 1 = 0 x^2 = 1 x = 1, -1

OpenStudy (anonymous):

ohhh well aren't I dumb/// sorry!! so x would be -1 and 1 !! AAA YES

ganeshie8 (ganeshie8):

You may continue your previous work like below : To find the x-intercept of C(x), we start by setting the equation ( C(x) ) to be equal to zero and then we solve for x. This can also be applied to any polynomial to find the x-intercept. For example C(x) = x^3-x. Setting this equal to 0 and solving gives the x intercepts x=0,-1,1. So the graph cuts the x axis at x=0,-1,1. A rough graph can be sketched by using x intercepts and the end bahavior. The degree of C(x) is odd and the leading coefficient is positive so the graph goes DOWN and UP on ends.

OpenStudy (anonymous):

so then with that I what is the second step?? @ganeshie8

OpenStudy (anonymous):

you are the best person ever omg thaNK YOU SO MUCH AAAAAA @ganeshie8 One more question and you won't hear from me again!! I'll tag you in the next one if that's okay with you?

ganeshie8 (ganeshie8):

I'll try.. .yw :)